OpenStudy (anonymous):

Is the equation true, false, or open? Explain. 9p+8=10p+7

OpenStudy (anonymous):

It's true because if you subtract 7 from both sides, you will 9p+1=10p Then, when you subtract 9p from both sides, you will have 1=1p so it's true

OpenStudy (anonymous):

It is open because true or false statements cannot have variables.

OpenStudy (anonymous):

Okay, Kell1170, is Math>philosphy right in any way?

OpenStudy (anonymous):

Well, if you divide by 1 on each side you have 1=p, which works, because when you plug it in, you have 9(1)+8 = 10(1)+7 So, 17=17.

OpenStudy (anonymous):

17 = 17 is a true sentence. 9p+8=10p+7 is neither true nor false.

OpenStudy (anonymous):

okay I just finished working the problem out, you are correct, so it appears! GO Girl!!!!!! Thank you sooo much!

OpenStudy (anonymous):

It asked if the equation was true or false, which you can't determine until you work it out.

OpenStudy (anonymous):

That's why it would have been an open statement, because it's undetermined until you solve for the variable

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

exactly, i just was confused as to where the heck to begin, I am soo greatful for you right now!

OpenStudy (anonymous):

Math>philosophy, if you went that route, every question would be open, which I highly doubt is the point of the problems. So, you must solve for the variable.

OpenStudy (anonymous):

Every question is open..I thought the point of the problem was determine logic/sets